OpenStudy (dinmosiaren):

Let f(x) = x2 + 6 and g(x) = . Find (g ° f)(–7).

OpenStudy (anonymous):

if g(x) need to find

OpenStudy (dinmosiaren):

x+8/x

OpenStudy (anonymous):

first subtitute your value g(f(x)) which mean put the function of f(x) inside the g(x) function

OpenStudy (anonymous):

do you get it.. for example f(x)=x+2 g(x)=x^2+3 g(f(x)) =g(x+2) =(x+2)^2+3________ here subtitute your f(x) inside your g(x)

OpenStudy (dinmosiaren):

okay and what do you get

OpenStudy (anonymous):

if g(x)=(x+8)/x or x+(8/x)

OpenStudy (dinmosiaren):

my answer has some fractions... -55/7 384/7 295/49 63/55

OpenStudy (anonymous):

(x+8)/x=63/55

OpenStudy (dinmosiaren):

thank you